Find the x and y intercepts and graph the equation.
-2x+7y=-7
-2x+7(0)=-7
-2x+0=-7
-2x=-7
x=7/2 ( 3 1/2, 0)
2(0)=7y=-7
0+7y=-7
7y=-7
y=0

I'm I on the right track?

Answer by mananth(16949)

-2x+7y=-7
-2x+7(0)=-7
-2x+0=-7
-2x=-7
x=7/2 ( 3 1/2, 0)
-2(0)+7y=-7
0+7y=-7
7y=-7
/7
y=-1 (0 , -1)